THERE ARE 850 DOUGLAS FIR AND HEMLOCK TREES IN A SECTION OF FOREST BOUGHT BY SAWZ LOGGING CO. THE COMPANY PAID AN AVERAGE OF $300 FOR EACH DOUGLAS FIR AND $225 FOR EACH PONDEROSA PINE. IF THE COMPANY PAID $217,500 FOR THE TREES, HOW MANY OF EACH KIND DID THE COMPANY BUY?
Answer by checkley71(8403)

D+P=850 OR P=850-D
300D+225(850-D)=217,500
300D+191,250-225D=217,500
75D=217,500-191,250
75D=26,250
D=26,250/75
D=350 DOUGLAS FIRS.
850-350=500 HEMLOCK/DOUGLAS FIRS.
PROOF
300*350=500*225=217,500
105,000+112,500=217500
217,500=217,500